40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 164     Entry time: Tue Jun 15 15:48:29 2010
Author: Frank 
Type: Computing 
Category: DAQ 
Subject: DAQ "no sync" problem fixed 

fixed the "no sync" problem we had since about three weeks. It was the 100pin flat ribbon cable from the timing adapter card to the A/D card. It looks ok but has problems at one of the crimped connections. If you touch it you can toggle the connection from "OK" to "not OK" via some states in between like "OK for 5 seconds". So i replaced that cable... Now DAQ is running with 4 sensors per chamber and one temp sensor for room temp measurements. will post all the channel names later...

ELOG V3.1.3-